OpenStudy (anonymous):

In how many ways can a gymnastics team of 4 be chosen from 9 gymnasts?

OpenStudy (anonymous):

9c4 Combination problem..!!

OpenStudy (ash2326):

@Chero did you understand?

OpenStudy (anonymous):

Im working it, hold one...

OpenStudy (ash2326):

sure :)

OpenStudy (anonymous):

not getting the right answer.

OpenStudy (anonymous):

my choices are: 36 126 3,024

OpenStudy (ash2326):

Would you share your work here?

OpenStudy (anonymous):

is it 3024??

Parth (parthkohli):

\( \color{Black}{\Rightarrow 9C4 = \Large {9! \over 4! 5! }}\)

OpenStudy (ash2326):

nope, you are missing something \[^AC_B=\frac{A!}{B!\times (A-B)!}\]

Parth (parthkohli):

Do you understand what factorials are?

OpenStudy (anonymous):

9!/(5! x4!) Now solve..:)

OpenStudy (anonymous):

braking it down?

Parth (parthkohli):

Yes. \(9! = 9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1 \)

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

Ans = (6x7x8x9) / (2x3X4)

OpenStudy (anonymous):

thank you!

Parth (parthkohli):

You're welcome :)